In a simple random sample of 1000 people, 261 are left-handed. Find a 95% confidence interval for the true proportion of left-handed people in the entire population
Answer by Boreal(15235)

the half interval is z(0.975)*sqrt(.261*.739/1000)=1.96*0.0139=0.0272
the interval is 0.261+/-0.0272
(0.234, 0.288)